OpenStudy (theeric):

Hi! This is a finance-based problem, which might be my difficulty. I have the answers from the book. Here is the prompt: "Suppose that a sum \(S_0\) is invested at an annual rate of return \(r\) and compounded continuously. (a) Find the time \(T\) required for the original sum to double in value as a function of \(r\). (b) Determine \(T\) if \(r=7\%\). (c) Find the return rate that must be achieved if the initial investment is to double in \(8\) years.

OpenStudy (theeric):

I think I should start out by finding \(S'(t)\), but that's where I think I'm wrong. It just comes from understanding the situation, I know.. I was thinking \(S'=S\ r\). Is that correct?

OpenStudy (theeric):

Oh, answers: (a) \(\ln(2)\div r\) year (b) \(9.90\) years (c) \(8.66\%\)

OpenStudy (anonymous):

Is it something like this?\[ 2S_0=S_0(1+r)^T \]

OpenStudy (anonymous):

Then solve for \(T\) first the first problem.

OpenStudy (theeric):

That doesn't look familiar... I used separation of variables to get \(S=e^{r\ t}\ C\), and then I was going to treat it like an initial value problem where \(S(0)=S_0\)

OpenStudy (anonymous):

Oh it's compounded continuously, I see.

OpenStudy (anonymous):

So you used a differential equation to solve this?

OpenStudy (theeric):

Yeah... I think that's what the class needs me to do! So, I need to think about what the rate is. Is it just the the \(S\) times the \(r\)? I was less sure when I asked... I think I'll see if I can get to a correct answer...

OpenStudy (anonymous):

What is your original differential equation?

OpenStudy (theeric):

I was thinking to use \(S'=S\ r\), and then from there I got to \(S=e^{rt}C\). I'm sorry, I'm doing a couple things at once, now. Don't worry about getting back to me - I'm only half here!

OpenStudy (anonymous):

Okay then your answer for (a) is correct.

OpenStudy (accessdenied):

Have you had luck with this problem so far? I think what you've had so far seems like it's on the right track, just curious. :)

OpenStudy (theeric):

Hi! I'm actually confused! The answers are from the book. I typed up my work here, but I hit a button and it all went away.. So I'm typing it up in notepad and will paste it here in a moment. Thanks for checking!

OpenStudy (accessdenied):

Alright, and yes that's a good idea. I do that from time to time as well in case of emergencies or when it feels laggy typing LaTeX stuff.

OpenStudy (theeric):

Then I did \(S(0)=S_0=e^{r\ (0)}C=C\implies S_0=C\) Then \(S=e^{r\ t}S_0\). I just realized I did some work but it doesn't pertain to the question!

OpenStudy (theeric):

Oh wait, maybe I was on the right track! One second!

OpenStudy (theeric):

Now I try to solve for \(T\). \(S=e^{r\ T}S_0\\\implies\dfrac{S}{S_0}=e^{r\ T}\\\implies\ln\left(\dfrac{S}{S_0}\right)=\ln(e^{r\ T})=r\ T\\\implies T=\ln\left(\dfrac{S}{S_0}\right)\ r^{-1}\) \(S=2S_0\\\implies T=\ln\left(\dfrac{2S_0}{S_0}\right)\ r^{-1}=\ln(2)\ r^{-1}\) Mission one, success! Thanks for sticking around to see my accomplishment! :)

OpenStudy (accessdenied):

Yes! You are welcome! The other two should follow up pretty nicely from your new equation now. :)

OpenStudy (theeric):

Is part (b) a follow-up to part (a)? Should I leave \(S=S_0\), do you think?

OpenStudy (theeric):

\(S=2S_0\), I mean.

OpenStudy (accessdenied):

yeah, I think they want you to simply apply your new equation. I was able to get the correct answers doing so.

OpenStudy (theeric):

(b) \(T=\dfrac{\ln(2)}{.07}\approx 9.90\) Thank you very much!

OpenStudy (theeric):

Thank you for checking the answers before I got to them, I mean! :D

OpenStudy (accessdenied):

Yep! I am glad to have been helpful! :)

OpenStudy (theeric):

:)

OpenStudy (theeric):

At this point, I found the answers! Thank you everyone! Just for completion, I'll show my work to get (c). Maybe this can help someone else, if they search for it or something. Or maybe if someone wants to use it as a reference! \(r=\ln(2)\ t^{-1}\) and \(t=8\) so \(r=\dfrac{\ln(2)}{8}=.08664...\approx.0866=8.66\%\) \(\huge\color{#44BB33}{\large\ \ o\ o\ \\\smile}\)